5-15 Present Value of an Annuity Find the present values of these ordinary annuities. Discounting occurs once a year.

  1. $600 per year for 12 years at 8%

  2. $300 per year for 6 years at 4%

  3. $500 per year for 6 years at 0%

  4. Rework parts a, b, and c assuming they are annuities due.
Answer #1

Ordinary annuity Annuity due 1 $ 4,521.65 $ 4,883.38 2 $ 1,572.64 $ 1,635.55 3 $ 3,000.00 $ 3,000.00

Ordinary annuity Annuity due 1 =PV(8%,12,-600) =PV(8%,12,-600,,1) 2 =PV(4%,6,-300) =PV(4%,6,-300,,1) 3 =PV(0%,6,-500) =PV(0%,